3. (x + y − 1)dx + (2x + 2y + 1)dy = 0


1
Expert's answer
2021-05-10T13:04:43-0400

We have given the differential equation,


"(x+y-1)dx+(2x+2y+1)dy = 0"


"\dfrac{dy}{dx} = - \dfrac{x+y-1}{2(x+y)+1} ----(1)"


Substituting "x+y = t"


then, "1 + \dfrac{dy}{dx} = \dfrac{dt}{dx}"


"\dfrac{dy}{dx} = \dfrac{dt}{dx} - 1"


Putting value in equation(1),


"\dfrac{dt}{dx} - 1 = \dfrac{1-t}{2t+1}"


"\dfrac{dt}{dx} = \dfrac{t+1}{t+\dfrac{1}{2}}"


Separating terms and integrating both sides,


"\int \dfrac{t}{t+1}dt + \dfrac{1}{2} \int \dfrac{dt}{t+1} = \int dx"


"t- log(t+1) + \dfrac{1}{2}log(1+t) = x + C"


"x+y-log(x+y+1)+ \dfrac{1}{2}log(1+t) = x+C"